# Localized Gaussian Splatting Editing with Contextual Awareness

### v1

@@ Abstract
Recent text-guided generation of individual 3D objects has achieved great success using diffusion priors. However, when directly applied to scene editing tasks (e.g., object replacement, object insertion), these methods lack the context of global information, leading to inconsistencies in illumination or unrealistic occlusion. To bridge the gap, we introduce an illumination-aware 3D scene editing pipeline for 3D Gaussian Splatting (3DGS) representation. Our key observation is that inpainting by the state-of-the-art conditional 2D diffusion model [57] excels at filling missing regions while considering the context. In this paper, we introduce the success of 2D diffusion inpainting to 3D scene object replacement. Specifically, our approach automatically picks the representative views that best express the global illumination, and these views are inpainted using the state-of-the-art 2D diffusion model. We further developed a coarse-to-fine optimization pipeline that takes the inpainted representative views as input and converts them into 3D Gaussian splatting to achieve scene editing. To acquire an ideal inpainted image, we introduce an Anchor View Proposal (AVP) algorithm to find a single view that best represents the scene illumination in target region. In the coarse-to-fine 3D lifting component, we first achieve image-to-3D lifting given an ideal inpainted view. In the fine step of Texture Enhancement, we introduce a novel Depth-guided Inpainting Score Distillation Sampling (DI-SDS), which enhances geometry and texture details with the inpainting diffusion prior, beyond the scope of the 3D-aware diffusion prior knowledge in the first coarse step. DI-SDS not only provides fine-grained texture enhancement, but also urges optimization to respect scene lighting. Our approach efficiently achieves local editing with global illumination consistency and we demonstrate robustness of our method by evaluating editing in real scenes containing explicit highlight and shadows, and compare against the state-of-the-art text-to-3D editing methods.
@@ 2 Related Works
The rising popularity of implicit 3D representations, particularly radiance fields [34, 18], has increased interest in 3D scene geometry and texture editing [10, 36, 26, 50, 29, 1, 59]. Advances in this field utilizes pre-trained vision, language and generative models [47, 4, 41, 23, 42], and achieves either direct modifications of entire scene  [51, 14, 9] or intricate manipulations and compositions of scene elements [1, 21, 12, 52, 33, 59].
Given a textual editing instruction, NeRF-Art [51] streamlines the modification of simple scenes by combining a foundational NeRF model with a style-oriented NeRF steered by CLIP priors. To progressively edit a NeRF scene, Instruct-NeRF2NeRF [14] alternatively updates training multi-views by Instruct-Pix2Pix [3] and then finetunes NeRF scene with updated training views. Later, to address the issue of multi-view inconsistent editing, ViCA-NeRF [9] enhances this process by integrating multi-view awareness with a blending mechanism. For more intricate scene alterations, SINE [1] advances image-guided 3D semantic editing by using an editing field informed by geometry and semantic vision priors [8, 4] to infuse detailed geometric and textural shifts. Meanwhile, some 3D editing works [48, 33, 59, 46, 45] focus on precise local changes with 3D segmentation and semantic editing. Within a target region in a NeRF scene, [48] achieves both geometry and texture alterations, and [46, 45] support object addition. DreamEditor [59] polishes local editing with a mesh-based field distilling from NeRF that can edit specific explicit segments within the 3D landscape. Concurrent works [11, 6] employ 3DGS for swift text-guided scene editing. Compared to these editing works, we resolve the problem of object generation and replacement with dramatic geometry changes and background alignment.
@@ 4.1 Anchor View Proposal
Existing multi-view diffusion models [28, 56] demonstrate the ability to generate consistent illumination across multi-views given the baked-in lighting in input single-view image. Inspired by this observation, we aim to identify an anchor viewpoint around the bounding box that includes the strongest illumination cues, such as shadow and highlight. To this end, we propose an Anchor View Proposal (AVP) algorithm to reliably select a well-conditioned view from rendering of $N_{anc}$ surrounding viewpoints looking at the center of bounding box. The ideal proposed view should contain the most significant difference in illumination between the left-right, top-bottom, or diagonal halves of the rendered images (see evaluation in Sec. 5.2 and Sec. 5.4). This is achieved by rotating each rendered image by a predefined set of angles and selecting the rotation that places majority of brighter pixels to the left. For robust lighting estimation from images, we convert rendered RGB images to HSV color space. This design choice is enlightened by the fact that V is the distance along the axis that turns black into non-black color, also known as “brightness”. Therefore, computation of V-channel image estimates amount of light that lit the surface. In scenarios where the majority of illumination is predominantly cast in the top-bottom direction, our algorithm tends to output an arbitrary view. This behavior aligns with our requirement, as we sample viewpoints in azimuth and any view is considered equally suitable. Pseudocode of AVP algorithm is provided in the supplementary material.
@@ 5.3 Comparison
Table 1: CLIP Score $\uparrow$ of compared cases in Fig. 6.
Vica-NeRF GaussianEditor Ours
Teddy Bear 0.5820 0.7461 0.7817
Buddha Statue 0.5596 0.7578 0.7725
Pikachu 0.5898 0.7939 0.8247
Astronaut 0.5518 0.6719 0.6748
Dinosaur 0.6777 0.7627 0.7354
Sculpture 0.5063 0.7275 0.7104
@@
We report CLIP Score  [15] for quantitative comparisons among baselines in Tab. 1. Specifically, we leverage open-sourced ViT-B/32 model of CLIP. We use foreground image inside mask to filter out influence of background when CLIP evaluates text description. We formulate text prompt by filling sentence “A photo of {}”, where text prompt in {} can be found in Fig. 6. Result shows that our generated objects are more aligned with input text prompt. In addition, we conducted a user study on the results and collected statistics of user’s preference. In the questionnaire, we put outputs of different methods in random order and ask participants which one is the most consistent with the text prompt and looks the most realistic within the scene. Statistics show that 0% users prefer Vica-NeRF, 29.49% prefer GaussianEditor and 70.51% prefer ours.
@@ 6 Conclusion
In this paper, we proposed an end-to-end pipeline to address text-guided localized editing of 3D scenes, focusing on generation of objects that naturally fit into the scene with illumination awareness. We observed that pre-trained 2D inpainting diffusion model is able to generate objects with texture consistent with scene illumination. Inspired by this, we proposed an Anchor View Proposal algorithm to find a viewpoint, under which we can obtain a conditioning image that contains the most contextual lighting information. We then proposed context-aware coarse-to-fine 3D generation pipeline that faithfully lifts conditioning image into 3D via a coarse step, and then adds details into geometry and texture via a Texture Enhancement step. To resolve the bottleneck that previous SDS does not support optimization with rendering background, we derived DI-SDS from depth inpainting ControlNet to supervise the Texture Enhancement. Experiments demonstrate that our pipeline successfully distills contextual illumination from the input scene.
@@ Acknowledgement
This research is sponsored by the U.S. Army under contract number W911NF-14-D-0005, Cooperative Agreement number W911NF-20-2-0053 and by the Office of Naval Research under contract number N00014-21-S-SN03. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army, Navy, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.

### v2

@@ Abstract
Recent advancements in text-guided 3D object generation using diffusion priors struggle with illumination inconsistencies when applied to scene editing tasks like object replacement or insertion. To address this, we propose an illumination-aware 3D scene editing pipeline for 3D Gaussian Splatting (3DGS). Our method leverages state-of-the-art 2D diffusion inpainting [62] to handle global illumination context effectively. Specifically, we identify representative anchor views that capture scene-wide illumination, inpaint them using 2D diffusion models, and integrate the results into a coarse-to-fine 3DGS optimization process. In the fine step, we introduce Depth-guided Inpainting Score Distillation Sampling (DI-SDS) to refine geometry and texture details, capitalizing on the diversity of 2D priors. Our approach achieves locally precise edits with globally consistent illumination, demonstrating robustness in real scenes with highlights and shadows. Comparisons show superior results over state-of-the-art text-to-3D editing methods. Project page: https://corneliushsiao.github.io/GSLE.html.
Abstract
In this supplementary material, we first detail the implementation specifics in Sec. 6.1, including the pseudocode of our Anchor View Proposal algorithm, the training configurations, and the hardware setup and training speed. Secondly, we supplement additional qualitative results in diverse scenes in Sec. 6.5. Thirdly, we present a user study demonstrating the superiority of our method compared to baseline approaches (Sec. 6.6). Fourthly, we showcase our graphical user interface in an example generation task (Sec. 6.7). Lastly, we discuss on the limitations and potential future directions of our research (Sec. 6.8). In addition to this document, we have attached several 360-degree videos and a live demo of our GUI usage.
@@ 2 Related Works
The rising popularity of implicit 3D representations, particularly radiance fields [36, 18], has increased interest in 3D scene geometry and texture editing [11, 38, 27, 54, 30, 2, 65]. Advances in this field utilizes pre-trained vision, language and generative models [49, 5, 43, 23, 44], and achieves either direct modifications of entire scene  [55, 15, 10] or intricate manipulations and compositions of scene elements [2, 21, 13, 56, 35, 65]. Given a textual editing instruction, NeRF-Art [55] streamlines the modification of simple scenes by combining a foundational NeRF model with a style-oriented NeRF steered by CLIP priors. To progressively edit a NeRF scene, Instruct-NeRF2NeRF [15] alternatively updates training multi-views by Instruct-Pix2Pix [4] and then finetunes NeRF scene with updated training views. Later, to address the issue of multi-view inconsistent editing, ViCA-NeRF [10] enhances this process by integrating multi-view awareness with a blending mechanism. For more intricate scene alterations, SINE [2] advances image-guided 3D semantic editing by using an editing field informed by geometry and semantic vision priors [9, 5] to infuse detailed geometric and textural shifts. Meanwhile, some 3D editing works [50, 35, 65, 48, 47] focus on precise local changes with 3D segmentation and semantic editing. Within a target region in a NeRF scene, [50] achieves both geometry and texture alterations, and [48, 47] support object addition. DreamEditor [65] polishes local editing with a mesh-based field distilling from NeRF that can edit specific explicit segments within the 3D landscape. Concurrent works [12, 7] employ 3DGS for swift text-guided scene editing. Compared to these editing works, we resolve the problem of object generation and replacement with dramatic geometry changes and background alignment.
@@ 4.1 Anchor View Proposal
Existing multi-view diffusion models [29, 60] demonstrate the ability to generate consistent illumination across multi-views given the baked-in lighting in input single-view image. Inspired by this observation, we aim to identify an anchor viewpoint around the bounding box that includes the strongest illumination cues, such as shadow and highlight. To this end, we propose an Anchor View Proposal (AVP) algorithm to reliably select a well-conditioned view from rendering of $N_{anc}$ surrounding viewpoints looking at the center of bounding box. Practically, we define a camera trajectory to render around the object and ensure capture of environmental illumination and no occlusion. The step may need certain amount of manual tuning such as distance to the object and elevation angle if the object in contact with a surface such as a table. The ideal proposed view should contain the most significant difference in illumination between the left-right, top-bottom, or diagonal halves of the rendered images (see evaluation in Sec. 5.2 and Sec. 5.4). This is achieved by rotating each rendered image by a predefined set of Azimuth angles and selecting the rotation that places majority of brighter pixels to the left. For robust lighting estimation from images, we convert rendered RGB images to HSV color space. This design choice is enlightened by the fact that V is the distance along the axis that turns black into non-black color, also known as “brightness”. Therefore, computation of V-channel image estimates amount of light that lit the surface. In scenarios where the majority of illumination is predominantly cast in the top-bottom direction, our algorithm tends to output an arbitrary view. This behavior aligns with our requirement, as we sample viewpoints in azimuth and any view is considered equally suitable. Pseudocode of AVP algorithm is provided in the supplementary material.
@@ 5.3 Comparison
We conducted a user study on the results and collected statistics of user’s preference. In the questionnaire, we put outputs of different methods in random order and ask participants which one is the most consistent with the text prompt and looks the most realistic within the scene. Statistics show that 0% users prefer Vica-NeRF, 29.49% prefer GaussianEditor and 70.51% prefer ours.
@@ 6 Conclusion
In this paper, we proposed an end-to-end pipeline to address text-guided localized editing of 3D scenes, focusing on generation of objects that naturally fit into the scene with illumination awareness. We observed that pre-trained 2D inpainting diffusion model is able to generate objects with texture consistent with scene illumination. Inspired by this, we proposed an Anchor View Proposal algorithm to find a viewpoint, under which we can obtain a conditioning image that contains the most contextual lighting information. We then proposed context-aware coarse-to-fine 3D generation pipeline that faithfully lifts conditioning image into 3D via a coarse step, and then adds details into geometry and texture via a Texture Enhancement step. To resolve the bottleneck that previous SDS does not support optimization with rendering background, we derived DI-SDS from depth inpainting ControlNet to supervise the Texture Enhancement. Experiments demonstrate that our pipeline successfully distills directional illumination from the input scene. Especially, we want to inspire the field to explore derivation and extension to SDS in conditioned generation and editing tasks.
@@ Acknowledgement
This research is sponsored by the U.S. Army under contract number W911NF-14-D-0005, Cooperative Agreement number W911NF-20-2-0053 and W911nf-25-2-0040 and by the Office of Naval Research under contract number N00014-21-S-SN03. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army, Navy, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Localized Gaussian Splatting Editing with Contextual Awareness
Supplementary Material
6.1 Implementation Details
6.2 Pseudocode of Anchor View Proposal Algorithm
We provide pseudocode for the proposed Anchor View Proposal (AVP) algorithm in Alg. 1. For the addConstrast function, we compute the minimum and maximum values in the V-channel image and map pixel values from the range [minimum, maximum] to [0, 1]. For the image plane rotation R, we select eight angles, each two views being 45 degrees apart in the 2D image plane.
6.3 Training Details
Resolution.
In the coarse step, we gradually increase the resolution of sampled random views from $64\times 64$ to $128\times 128$, and then to $256\times 256$ at iterations 0, 200, and 300, respectively. For the conditioning view, we increase the resolution from $128\times 128$ to $256\times 256$, and then to $512\times 512$ at the same iteration milestones. We found that the coarse step often converges around 3500 iterations, generating relatively fine detail. During the Texture Enhancement step, the resolution of both random views and the conditioning view is set to $512\times 512$.
Classifier-free Guidance Scale.
We use a guidance scale of 3.0 in the coarse step and 100 for DI-SDS in the Texture Enhancement step.
Algorithm 1 Anchor View Proposal
Input: multi-view RGBA images $I$, leftIsBrighter flag $f_{l}$
Output: anchor view index $\hat{i}$
function anchorViewProposal(II, $f_{l}$)
  $V\leftarrow$ getValueChannel(RGBA2HSV(II)) $\triangleright$ Value Images
  $M\leftarrow$ getAlphaChannel(II) $\triangleright$ Mask
  $R\leftarrow[1\ldots 8]\times 45$ $\triangleright$ Image Plane Rotations
  $B\leftarrow\mathbf{0}$ $\triangleright$ Image Brightness Balance
  for all $v_{i}$, $m_{i}$ in $V$, $M$ do
   $S\leftarrow\mathbf{0}$
   $v_{i}[m_{i}]\leftarrow$ addContrast($v_{i}[m_{i}]$)
   for all $r$ in $R$ do
    $g_{i,r}\leftarrow$ rotateImage($v_{i}$, $r$)
    $m_{i,r}\leftarrow$ rotateImage($m_{i}$, $r$)
    $S[r]\leftarrow$ calculateBrightnessBalance($g_{i,r}$, $m_{i,r}$, $f_{l}$)
   end for
   $B[i]\leftarrow$ min(S1S-1)
  end for
  $B\leftarrow$ abs($B$)
  return $\hat{i}\leftarrow$ argmin(B)
end function
function calculateBrightnessBalance($v$, $m$, $f_{l}$)
  $h,w\leftarrow$ imageHeight($v$), imageWidth($v$)
  $I_{l},I_{r}\leftarrow$ leftHalf($v[m]$), rightHalf($v[m]$)
  $b_{l},b_{r}\leftarrow$ mean($I_{l}$), mean($I_{r}$) $\triangleright$ Brightness of Masked Half Pixels
  if $f_{l}$ is True then
   return $b_{l}/(b_{l}+b_{r})$
  else
   return $b_{r}/(b_{l}+b_{r})$
  end if
end function
Loss Weights.
In the coarse step, the weight of the RGB loss $\lambda_{rgb}$ and the weight of the mask loss $\lambda_{mask}$ both start from 0, increase to 1500 at iteration 1000, and stay at 1000 for the remaining iterations. The weight of the optimization loss by the 3D-aware diffusion prior $\lambda_{3D-SDS}$ remains constant at 1. During the Texture Enhancement step, $\lambda_{rgb}$ and $\lambda_{mask}$ are constantly 1500, and the weight of the DI-SDS loss $\lambda_{DI-SDS}$ is constantly 0.1.
Learning Rates.
The learning rate for the position of Gaussian splats starts from $1\times 10^{-3}$ and decreases to $10^{-5}$ with exponential decay from iteration 0 to 1000 in the coarse step, and from $10^{-4}$ to $10^{-5}$ from iteration 0 to 10000 in the Texture Enhancement step. The learning rate for the scaling of Gaussian splats is $5\times 10^{-3},5\times 10^{-4}]$ in the coarse step and $5\times 10^{-4},5\times 10^{-5}]$ in the Texture Enhancement step, with the same milestones as the position. The rotation, color, and alpha of Gaussian splats are constantly $5\times 10^{-3},10^{-2},10^{-2}$, respectively, in both steps.
6.4 Hardware Requirement & Speed
We train our pipeline on a single NVIDIA A40 GPU. The VRAM usage is approximately 20GB for the coarse step and 43GB for the Texture Enhancement step. Each iteration takes around 0.5 seconds for the coarse step and 7.8 seconds for the Texture Enhancement step, as per the settings detailed in Sec. 6.3. The total number of iterations depends on complexity in geometry and texture. Typically, we found the coarse step converges after 6500 iterations. While texture enhancement step highly depends on user’s preference, we usually stops this step after 2000 iterations incremented from the coarse step. We sample random views with a batch size of 6 for the coarse step and 4 for the Texture Enhancement step. We discovered that increasing the batch size significantly reduces color deviation in the Texture Enhancement step.
6.5 More Results
We provide more qualitative results in Fig. 8,  9 and  10. In the figures, we show the anchor view images of the original scenes with 3D bounding box for editing, inpainted anchor view images, text prompts for generation, and multi-view rendering results from our pipeline. In addition to LERF [19], MipNeRF360 [3] and our self-captured datasets, we use a commercially purchased synthetic scene (the first case in Fig. 8 and the third case in Fig. 10). For the synthetic scene, we rendered RGB images in $800\times 800$ resolution under 332 virtual camera views ($48.5$-degree field of view) around the scene to densely supervise 3DGS training. For our self-captured datasets, we use a Canon EOS-1D camera to capture around 400 photos of each scene and send the raw colored images to COLMAP for calibration.
6.6 User Study
Figure 7: User Study statistics
To further substantiate our claim that our method generates more photorealistic results than state-of-the-art methods – GaussianEditor [7] and Vica-NeRF [10], we conducted an anonymous user study. The statistics of users’ preferences are provided in Fig. 7.
We created a questionnaire featuring output examples from different methods shown in Fig.5 of the main paper. The order of the methods was randomized to mitigate potential bias. Participants were asked to choose the result that best aligned with the input text prompt and appeared the most photorealistic in each question comparing the three methods.
We also requested participants to share their age group (options were “‘below 18”, “18-25”, “26-35”, “36-45”, “46+”) and their familiarity with the fields of computer vision (CV) or computer graphics (CG) (options were “Yes”, “No”, “Maybe”). All 26 questionnaires were collected anonymously.
According to the results, no users preferred Vica-NeRF, 29.49% preferred GaussianEditor, and 70.51% preferred our method. Among the participants, 19.2% fell into the age group of 18-25, 76.9% into 26-35, and 3.8% did not provide their age. As for familiarity with CV or CG, 61.5% claimed familiarity, 26.9% claimed no familiarity, 7.7% were uncertain (“Maybe”), and 3.8% did not provide their familiarity level.
6.7 Graphical User Interface
We also package our pipeline into an application with graphical user interface as an extra contribution. An example of the interface can be found in Fig. 11. The system is built upon Viser [1], an open-source interactive 3D visualization library. The system supports original GSGen [8] and 3D Gaussian Splatting [18] rendering and live progress check of training. We also leave APIs easy to change in a separate configuration file. This allows users and researchers to conduct future experiments in both text-to-3D and image-to-3D generative tasks.
6.8 Discussion
6.9 Limitations
Our current method guarantees photorealism between the generated object and input scene, but it does not ensure physically accurate lighting. In scenes with complex lighting, a single conditioning anchor view often provides insufficient lighting information. Additionally, current text-guided generation schemes do not support shadow editing, which requires disentanglement and editing of background. Therefore, it is necessary to involve conventional inverse rendering and ray tracing similar to [61, 34, 51, 63, 53, 25]. Also, our pipeline may still require manual tuning in AVP step, where the collection of renderings for anchor view proposal needs to avoid occlusion and surface intersection. When the directional illumination is cast from top to bottom, the manual set elevation angle is crucial. It will be intuitive and straightforward to extend Azimuth rotation to further sample also elevation rotation, but this will also require manual tuning.
6.10 Future Works
Our current task involves object-centered replacement and editing. This process requires the editing region to be enclosed in a bounding box, separate from the other point clouds in the scene. As a result, detailed editing that interacts with existing scene objects is not yet possible. In such cases, we need a more precise segmentation and generation method to optimize both the generated point cloud and the scene point cloud, to ensure multi-view consistency.
Figure 8: More Results 1 of 3. We show original and inpainted anchor view in the first column, and multi-view rendering results of our method in the right three columns.
Figure 9: More Results 2 of 3. We show original and inpainted anchor view in the first column, and multi-view rendering results of our method in the right three columns.
Figure 10: More Results 3 of 3. We show original and inpainted anchor view in the first column, and multi-view rendering results of our method in the right three columns.
Figure 11: Graphical User Interface of our pipeline. We develop a GUI that includes Anchor View Proposal, coarse image-to-3D generation and texture enhancement steps. The interactive viewer is especially useful to visualize training procedure and results. In the lower left figure, we show rendering of depth images of the same scene besides rendering of RGBA frames.